OpenStudy (anonymous):

The equation of the circle with center (3, -2) and radius 7 is:

OpenStudy (sumi29):

@adrian1220 : Are you familiar with the general equation of a circle? If you are, then you can just substitute the data. Its simple.

OpenStudy (anonymous):

can you take me step by step then

OpenStudy (sumi29):

Sure. Ok, the general equation for a circle is: \[(x+a)^2 + (y+b)^2 = r^2\] where (a,b) is the center co-ordinate, and r is the radius. Now just put the data given into this equation.

OpenStudy (anonymous):

so it (x+3)^2+(y+2)^2=14